OpenStudy (anonymous):

Find the distance of (1,0,0) from the line r=t(12i -3j -4k)

OpenStudy (anonymous):

I shall take a numerical example because then I can explain both the principles and how the working out goes. Suppose the position vector of P from the origin O is OP = p = (2, -1, 2). Suppose that the vector equation of line L is r = (-1, 0, 7) + t (4, 1, -2). The shortest distance of P from L is given by the length of the perpendicular from P to L. Suppose this perpendicular meets L at H. Then we want to find the length of PH. There are two ways we can set about this. Method (1) Using the dot product Since H lies on L we can say that OH = h = (-1, 0, 7) + t (4, 1, -2) = (-1 + 4t, t, 7-2t) for some value of t which we need to find. Also, vector PH = -p + h = - (2, -1, 2) + (-1 + 4t, t, 7 - 2t) = (-3 + 4t, 1 + t, 5 - 2t). But PH is perpendicular to the direction vector (4, 1, -2) of line L. So the dot product of vector PH and (4, 1, -2) is zero. So (-3 + 4t, 1 + t, 5 - 2t).(4, 1, -2) = -12 + 16t + 1 + t - 10 + 4t = - 21 + 21t = 0 so t = 1. Therefore OH = h = (3, 1, 5) and vector PH = -p + h = (1, 2, 3). Its length |OH| is given by the square root of (1 + 4 + 9) = 3.74 to 2 sf. Method (2) Using the equation of a plane. Just as we found in method (1), we have h = (-1 + 4t, t, 7-2t). Now PH lies in a plane which is perpendicular to line L. So the direction vector (4, 1, -2) of L is perpendicular to this plane. Therefore it is a normal vector to the plane. Also, P lies in the plane. Using the equation of a plane of r.n = p.n, with r = (x, y, z), we get (x, y, z).(4, 1, -2) = (2, -1, 2).(4, 1, -2) so 4x + y - 2z = 3. But H also lies in this plane so we can say 4 (-1 + 4t) + t - 2 (7 - 2t) = 3 so 21t = 21 so t = 1 as before. hope this helps u!!

OpenStudy (anonymous):

This doesn't help me much...

OpenStudy (anonymous):

how

OpenStudy (anonymous):

I don't know how to use this...

OpenStudy (anonymous):

the answer is\[ \sqrt {69} \over 13 \]but I got 1/13

OpenStudy (anonymous):

??

OpenStudy (experimentx):

tan(alpha) = sqrt(1-cos2(alpha))/cos(alpha) = 5/13/12/13 = 5/12 5/12 = d/1 => d = 5/12 ???

OpenStudy (anonymous):

that's not the answer...?